# Directional Diffusion-Style Code Editing Pre-training

### v1

@@ VIII Conclusion
In this paper, we propose DivoT5, a directional diffusion-style pre-trained model designed to simulate the incremental nature of human code editing and effectively capture the characteristics of code evolution. Specifically, we first propose three pre-training tasks to make the model aware of fine-grained code editing information (i.e., the direction of code evolution) and enrich the context of the old code. Then, we design the fourth pre-training task utilizing intermediate states during the editing process to further reinforce the direction of code evolution. Experiments are conducted on two code-editing scenarios and one non-editing scenario using five datasets, demonstrating that DivoT5 achieves state-of-the-art performance on all tasks compared to fine-tuned similar or larger scale models (60M, 220M, 770M), and few-shot billion-scale models (6.7B, 8B, ChatGPT). In addition, our ablation experiments also confirm the contribution of each pre-training task to the final performance.

### v2

@@ VIII Conclusion
In this paper, we propose DivoT5, a directional diffusion-style pre-trained model designed to simulate the incremental nature of human code editing and effectively capture the characteristics of code evolution. Specifically, we first propose three pre-training tasks to make the model aware of fine-grained code editing information (i.e., the direction of code evolution) and enrich the context of the old code. Then, we design the fourth pre-training task utilizing intermediate states during the editing process to further reinforce the direction of code evolution. Experiments are conducted on two code-editing scenarios and one non-editing scenario using eight datasets, demonstrating that DivoT5 achieves state-of-the-art performance on most tasks compared to fine-tuned similar or larger scale models (60M, 220M, 770M, 6.7B), and few-shot billion-scale instruct models (6.7B, 8B, ChatGPT). In addition, our ablation experiments also confirm the contribution of each pre-training task to the final performance.
IX Acknowledgment
This work is sponsored by the National Key Research and Development Program of China under Grant No. 2023YFB4503803, the National Natural Science Foundation of China under Grant No. 62232003 and No. 62402482, and the CCF-Huawei Populus Grove Fund (Grant No. CCF-HuaweiSE202407).